Let f(x) = 3r +44.23 + 204r? + 288. - 3. Calculate the derivative f(x) = Calculate the second derivative f(x) Note interv
Answer #1

3 2 Given f«) = 324 +44+ 2043 +2883–3. differentiate f(x) with respect to a. fCod (3x++4422+ 204** +2888-3) dx = 3(42)+44(3|-6<3 <-41 213-5 interval | x-value f(x) = 192271322+406x+288 Comment -alal-6 x=-7 12 6-7)+1326-2)5 4081-74299 de creating -between x < x <- 5.120 The value Second derivative is Positive and also -9.32loo The value of second derivative is Positive C